OpenStudy (anonymous):

Is -19.78 a irrational or rational # ????

OpenStudy (anonymous):

Rational because it can be written as a fraction while a number like 3.8873542..... cannot be.

OpenStudy (anonymous):

Okay thank you! what about something like this though...? \[\sqrt{-3}/5 + 10.8\]

OpenStudy (anonymous):

Irrational or rational?

OpenStudy (austinl):

That is irrational because you would get imaginary numbers. Because you have a negative under the radical.

OpenStudy (anonymous):

So if its negative its always irrational???

OpenStudy (anonymous):

-0.8662866239 would be irrational?

OpenStudy (phi):

irrational numbers have an unending, non-repeating decimal digits For example, if you find the square root of 3 your calculator gives 1.732050808 but if your calculator printed more digits you would see \[ \sqrt{3} \approx 1.732050807568877293527446341505872366942805253810380628055806979451933016908800037081146186757248576 \] and it really goes on forever

OpenStudy (anonymous):

Okay, that helps a lot thank you!!!

OpenStudy (phi):

square root and cube roots are almost always irrational except for special numbers (perfect squares or perfect cubes) for example sqr(24) is irrational sqr(25) = 5 (because 25 is a perfect square= 5*5)

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

Im getting it a lot more now